OpenStudy (anonymous):

find the value of x if 6/x=3/4

OpenStudy (amistre64):

6/2 = 3 x/2 = 4 x=8

OpenStudy (amistre64):

or...6/3 = x/4 x=8

OpenStudy (amistre64):

or... 6(4)/3 = x 24/3 = x 8 = x

OpenStudy (anonymous):

that makes sence

OpenStudy (amistre64):

makes thrice sense :)